# Parametric Graph Representations in the Era of Foundation Models: A Survey and Position

### v1

@@ Abstract
Graphs have been widely used in the past decades of big data and AI to model comprehensive relational data. When analyzing a graph’s statistical properties, graph laws serve as essential tools for parameterizing its structure. Identifying meaningful graph laws can significantly enhance the effectiveness of various applications, such as graph generation and link prediction. Facing the large-scale foundation model developments nowadays, the study of graph laws reveals new research potential, e.g., providing multi-modal information for graph neural representation learning and breaking the domain inconsistency of different graph data. In this survey, we first review the previous study of graph laws from multiple perspectives, i.e., macroscope and microscope of graphs, low-order and high-order graphs, static and dynamic graphs, different observation spaces, and newly proposed graph parameters. After we review various real-world applications benefiting from the guidance of graph laws, we conclude the paper with current challenges and future research directions.
Figure 1: A Toy Example of Illustrating Graph Laws in the Real World: Different types of graphs follow some shared properties that can be used to parameterize their structure. The left figure depicts a social network, and the right figure represents flight routes between several major cities after COVID-19, where flights gradually recover to their pre-pandemic state. Both graphs exhibit two key properties: (1) Global Shrinking Law: As the graph evolves, the diameter decreases, indicating a more connected structure. (2) Local Triangle Closure Law: New links are more likely to close triangles. These behaviors are indicative of typical graph dynamics seen in social networks, transport systems, and other domains of graphs. These graph laws have the potential to be quantified and serve as fundamental principles for developing graph foundation models.
@@ 1 Introduction
In the era of big data and AI, graphs are popular data structures for modeling the complex relationships between entities. Also, graph-based research (e.g., graph mining and graph representation learning) provides the foundation for many real-world applications, such as recommender system [46, 22, 49, 51], social network analysis [35, 23], information retrieval [36, 45, 32], anomaly detection [63, 64, 17], natural language processing [50, 7, 65], computer vision [8, 27], AI4Science [59, 18, 15], etc.
@@
To model those real-world tasks within graphs, graph representations are indispensable middleware that provides the basis for specific and complex task-oriented computations. To be specific, graph representations can be decomposed into three aspects, (1) graph embedding (i.e., vector representation), (2) graph law (i.e., parametric representation), and (3) graph visualization (i.e., visual representation). First of all, graph representations can be in the form of embedding matrices, i.e., the graph topological information and attributes are encoded into matrices, which has been widely discussed and studied in the research community and usually can be referred to as graph representation learning[20]. Then, graphs can be also represented by plotting directly for a better human-understandable illustration. For example, one interesting research topic is how to plot the graph topological structures into the 2D space with less structure distortion. More interesting works can be referred to [16]. Last but not least, graphs can also be represented by a few key parameters such as Erdős-Rényi random graph $G(n,p)$ [12], where $n$ stands for the number of nodes in the graph, and $p$ stands for the independent edge connection probability in the graph. As shown in Figure 2, these three representations can have overlapping to mutually contribute to each other [16].
1.1 Motivation of this Paper
In the modern graph deep learning community, graph embedding (also referred to as graph representation learning) has attracted unprecedented research attention varying from unsupervised (e.g., Node2Vec [19]) to semi-supervised (e.g., GCN [28]), and also the novel neural architecture from transformer (e.g., GraphTransformer [13]). However, the graph law (also referred to as graph parametric representation) has great research potential, yet the corresponding research stays in a not fully exploited stage.
At the beginning, we would point out “What is Graph Law or Graph Parametric Representation?”. In general, it is referred to as finding some key parameters and their relations to describe the given graph, which description is expected to preserve the entire or part of the property of the given graph. According to the previous research [33, 34, DBLP:conf/kdd/FuFMTH22], the rigorous graph law (or graph parametric representation) relies two fundamental steps. The first step is to decide which parameters we use to describe the graph, e.g., the node degree. The second step is to observe the graph in a statistical manner and then determine the value of parameters, the distribution of parameters, the relation among parameters, etc. For example, the relation between the possibility of a newly-arrived node connecting to an old node (parameter 1) and the degree of that old node (parameter 2) is studied by maximum likelihood estimation (MLE) based on the observed real-world graph data [34].
Then, we discuss why studying the graph parametric representation is important in the graph research community nowadays with the following two concrete examples.
  • Foundation of Graph Foundation Models. Similar to building the foundation models in other modalities [61, 62, 3], Graph data integration is the base to support various graph AI developments by aggregating different domain graph data for extracting non-trivial, abundant, and useful knowledge. But its realization is very challenging, for example, unlike the independent and identically distributed (IID) data, the distribution of graph data from different domains can be quite different and not easy to be jointly leveraged, e.g., the graph size, the attribute dimension, and the attribute meaning of different domain graphs place a barrier for distilling a consensus intelligence [44]. Graph law acts like a potential trigger to break the inconsistency of different domains, such that different graphs can be described under the same statistical language. With a suite of powerful graph laws, the cross-domain graph (or subgraph) representation complexities can be reduced to several shareable parameters, akin to Erdős–Rényi graphs but accounting for heterogeneity and temporality, such that the large-scale foundation model training among various graph data can become promising.
  • Bring Graphs to Foundation Models. Whether the input graph statistical property gets well preserved during the representation learning process has become an increasingly interesting research question recently, and some preliminary theoretical studies [48] will be discussed in the following part of this paper. Further, take the concrete application scenarios as examples, in the climate domain, modeling the geolocation as graphs [31], graph neural networks (GNNs) have been deployed for the weather forecasting and obtain the outperformance. Although understood by machines, a follow-up question gets raised as to whether this GNN-encoded knowledge obeys the physical law or equation of climate, i.e., does the machine understand the climate as humans do? Thus, a possible solution is to find a way to model the graph law in terms of physical rules, try to encode the law along with the neural representation learning process, and see the decision variance to verify the hypothesis. In addition to natural science, graph law also has the potential to bring human activities into the neural representation learning process and boost task performance. For example, one recent study [25] shows that adding extra local topology information into the prompt can help large language models (LLMs) achieve textual node classification tasks with high accuracy. More discussion on graph laws and LLMs like [53, 21] can be found in Section 6.6.
Figure 3: Organization of Graph Law Introduction.
@@ 1.2 Organization of this Paper
Graph law is the study of investigating the statistical properties of graphs. In this survey, we introduce the graph laws studies in the macroscopic view and microscopic view, plus multiple angles like low-order and high-order connections, static and dynamic graphs, as shown in Table 1 and Figure 3.
  • Macroscopic and Microscopic Views. The macroscopic graph laws describe the graph properties in a global view, like how the total degree (or eigenvalues) distribution of the entire graph looks like [33]; while the microscopic laws try to focus on the individual behavior and investigate their behaviors as part in the entire graph [34].
  • Low-Order and High-Order Connections. Most graph laws are based on the node-level connections (i.e., low-order connections), while some graph law investigations are based on the group activities (high-order connections), i.e., motif in [40, 57], hyperedge in [10, 29], and simplex in [6, 9].
  • Static and Dynamic Graphs. Compared with static graphs, dynamic graphs allow the graph components like topology structure and node attributes to evolve over time. Correspondingly, some graph laws study how the graph parameters change over time and their temporal relations. Note that, in some graph research, the dynamics are created by the algorithms, like adding virtual nodes to preserve graph representations [1, 37]; this kind of research is beyond the discussion of this paper, and we focus on the graph law with natural time.
Moreover, in Section 4, we introduce different observation spaces and newly proposed parameters, before the corresponding laws are discovered.
Furthermore, in Section 5, we survey different real-world applications that would benefit from the guidance of graph laws, such as graph generation, link prediction, and natural language processing.
Finally, in Section 6, we conclude the survey with current challenges and 6 future directions for graph law research.
@@
Graphs Densification Law [33] Density degree $\alpha$ Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\in[1,2]$, $e(t)$ is # edges at $t$
@@
Shrinking Law [33] Effective diameter $d$ Macro Low Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
@@
Motif Differing Law(1) [40] Numbers of similar motifs $n$ Macro High Dynamic $n_{1}\neq n_{2}$ for different domains
@@
Motif Differing Law(2) [40] Motif occurring timestamp $t$ Macro High Dynamic $t_{1}\neq t_{2}$ for different motifs
@@
Egonet Differing Law [6] Features of Egonets $X$ Macro High Static $X_{1}\neq X_{2}$ for different domains
@@
Simplicial Closure Law [6] Simplicial closure probability $p$ Macro High Static $p$ increases with additional edges or tie strength
@@
Spectral Power Law(1) [14] Degree, SVD, eigen distributions Macro High Static These distributions usually follow power-law
@@
Spectral Power Law(2) [14] Degree, SVD, eigen distributions Macro High Static If one follow power-law, usually others follow
@@
Edge Attachment Law(1) [34] Node degree $d$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)\propto d$ for node with degree $d$
@@
Edge Attachment Law(2) [34] Node age $a(u)$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)$ seems to be non-decreasing with $a(u)$
@@
Triangle Closure Law(1) [24] Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{3}\Rightarrow$ unlikely $e_{1}/e_{2}$ will be weakened
@@
Triangle Closure Law(2) [24] Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{1}/e_{2}\Rightarrow$ unlikely they will be weakened
@@
Local Closure Law [54] Local closure coefficient $H(u)$ Micro Low Static Please refer to Section 4 for details
@@
Spectral Density Law [11] Density of states $\mu(\lambda)$ Macro High Static Please refer to Section 4 for details
@@
Motif Activity Law(1) [57] Motif type Micro High Dynamic Motifs do not transit from one type to another
@@
Motif Activity Law(2) [57] Motif re-appear rate Micro High Dynamic Motifs re-appear with configured rates
@@
Hypergraphs Degree Distribution Law [10] Node degree, edge link probability Macro High Dynamic High-degree nodes are likely to form new links
@@
SVD Distribution Law [10] Singular value distribution Macro High Static Singular value distribution usually heavy-tailed
@@
Diminishing Overlaps [29] density of interactions $DoI(\mathcal{H}(t))$ Macro High Dynamic Overall hyperedge overlaps decrease over time
@@
Densification Law [29] Density degree $\alpha$ Macro High Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$, $e(t)$ is # hyperedges at $t$
@@
Shrinking Law [29] Hypergraph effective diameter $d$ Macro High Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
@@
Edge Interacting Law [9] Edge interacting rate Micro High Dynamic Temporally adjacent interactions highly similar
@@
Heterographs Densification Law [47] Density degree $\alpha$, # meta-path Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$ for some meta-path
@@
Non-densification Law [47] Density degree $\alpha$, # meta-path Macro Low Dynamic Maybe, for some meta-path, $e(t)\not\propto n(t)^{\alpha}$
@@ 2 Macroscopic Graph Laws
Several classical theories model the growth of graphs, for example, Barabasi-Albert model [4, 5] assumes that the graphs follow the uniform growth pattern in terms of the number of nodes, and the Bass model [38] and the Susceptible-Infected model [2] follow the Sigmoid growth (more random graph models can be founded in [12]). However, these pre-defined graph growths have been tested that they could not handle the complex real-world network growth patterns very well [30, 56]. To this end, researchers begin to fit the graph growth on real-world networks directly, to discover graph laws.
@@ 2.1 Low-Order Macroscopic Laws
Based on fitting nine real-world temporal graphs from four different domains, the authors in [33] found two temporal graph laws, called (1) Densification Laws and (2) Shrinking Diameters, respectively. First, the densification law states as follows.
@@
where $e(t)$ denotes the number of edges at time $t$, $n(t)$ denotes the number of nodes at time $t$, $\alpha\in[1,2]$ is an exponent representing the density degree. The second law, shrinking diameters, states that the effective diameter is decreasing as the network grows, in most cases. Here, the diameter means the node-pair shortest distance, and the effective diameter of the graph means the minimum distance $d$ such that approximately 90% of all connected pairs are reachable by a path of length at most $d$. Later, in [56], the densification law gets in-depth confirmed on four different real social networks, the research shows that the number of nodes and number of edges both grown exponentially with time, i.e., following the power-law distribution.

### v2

# What Do LLMs Need to Understand Graphs: A Survey of Parametric Representation of Graphs

@@ Abstract
Graphs, as a relational data structure, have been widely used for various application scenarios, like molecule design and recommender systems. Recently, large language models (LLMs) are reorganizing in the AI community for their expected reasoning and inference abilities. Making LLMs understand graph-based relational data has great potential, including but not limited to (1) distillate external knowledge base for eliminating hallucination and breaking the context window limit for LLMs’ inference during the retrieval augmentation generation process; (2) taking graph data as the input and directly solve the graph-based research tasks like protein design and drug discovery. However, inputting the entire graph data to LLMs is not practical due to its complex topological structure, data size, and the lack of effective and efficient semantic graph representations. A natural question arises: Is there a kind of graph representation that can be described by natural language for LLM’s understanding and is also easy to require to serve as the raw input for LLMs? Based on statistical computation, graph laws pre-defines a set of parameters (e.g., degree, time, diameter) and identifies their relationships and values by observing the topological distribution of plenty of real-world graph data. We believe this kind of parametric representation of graphs, graph laws, can be a solution for making LLMs understand graph data as the input. In this survey, we first review the previous study of graph laws from multiple perspectives, i.e., macroscope and microscope of graphs, low-order and high-order graphs, static and dynamic graphs, different observation spaces, and newly proposed graph parameters. After we review various real-world applications benefiting from the guidance of graph laws, we conclude the paper with current challenges and future research directions.
@@ 1 Introduction
1.1 Motivation of this Paper
Graphs serve as a fundamental relational data structure and are extensively utilized in a wide range of application scenarios, including molecule design, social network analysis, and recommender systems Zhou et al. (2020). Their ability to represent complex interconnections among entities makes them indispensable in modeling real-world relationships. However, despite their widespread use, integrating graph-based data input with large language models (LLMs) remains a challenging problem, as shown in Figure 2 with a ChatGPT 4o Hurst et al. (2024) case study.
Recently, LLMs have revolutionized the AI community with their remarkable reasoning and inference capabilities Touvron et al. (2023); Guo et al. (2025), and those models have demonstrated significant potential in various tasks, such as natural language understanding, machine translation, and knowledge extraction. Given the growing importance of LLMs, enabling them to comprehend and process graph-based relational data could open new frontiers in artificial intelligence research and applications. This integration holds immense potential for enhancing LLMs in multiple ways, including but not limited to:
  • Knowledge Distillation for LLMs: Graph-based external knowledge bases can provide crucial insights, mitigating issues such as hallucinations in LLM-generated responses and overcoming the limitations imposed by fixed context windows. By incorporating structured graph data, LLMs can improve retrieval-augmented generation (RAG) techniques and enhance inference accuracy Edge et al. (2024); He et al. (2024b).
  • Direct Graph-Based Problem Solving: Many research domains, such as protein design and drug discovery, inherently rely on graph-based data representations Liang et al. (2023); Wang et al. . Equipping LLMs with the capability to understand and manipulate graph structures could significantly advance research in these fields by enabling direct problem-solving approaches.
Despite the clear advantages of incorporating graph data into LLMs, several challenges hinder this integration. The primary obstacles include (1) the complexity of graph topologies, (2) the size of graph datasets, and (3) the absence of effective semantic representations of graphs that LLMs can process efficiently. Unlike textual data, which LLMs are inherently designed to understand, graphs lack a straightforward natural language representation. This leads to a fundamental research question: Is there a form of graph representation that is both interpretable in natural language for LLMs and compact enough to serve as a viable input format?
A promising solution lies in the concept of graph laws. Graph laws refer to statistical principles that define relationships between key structural parameters of graphs, such as degree, clustering coefficients, diameter, and time. Hence, a graph can be represented by a few parameters to reflect its properties well. Correspondingly, the parameters’ formal mathematical relations and values are captured by graph law researchers by analyzing real-world, large-scale graph data distributions. By encoding graph properties through predefined sets of parameters, graph laws offer a way to translate complex graph topologies into a form that LLMs can potentially comprehend. For example, according to the previous research Leskovec et al. (2005, 2008), the relation between the possibility of a newly-arrived node connecting to an old node (parameter #1) and the degree of that old node (parameter #2) is studied by maximum likelihood estimation (MLE) based on the observed real-world graph data Leskovec et al. (2008).
1.2 Background of Graph Representations
To model real-world tasks within graphs, graph representations are indispensable middleware that provides the basis for specific and complex task-oriented computations. To be specific, graph representations can be decomposed into three aspects as shown in Figure 1, (1) graph embedding (i.e., vector representation), (2) graph law (i.e., parametric representation), and (3) graph visualization (i.e., visual representation).
@@ 1.2 Background of Graph Representations
Figure 2: A Case Study of ChatGPT 4o’s Explaining about Why it Needs Graph Laws
@@
First of all, graph representations can be in the form of embedding matrices, i.e., the graph topological information and attributes are encoded into matrices, which has been widely discussed and studied in the research community and usually can be referred to as graph representation learningHamilton (2020). Then, graphs can also be represented by plotting directly for a better human-understandable illustration. For example, one interesting research topic is how to plot the graph topological structures into the 2D space with less structure distortion. More interesting works can be referred to Fu and He (2022). Last but not least, graphs can also be represented by a few key parameters such as Erdős-Rényi random graph $G(n,p)$ Drobyshevskiy and Turdakov (2020), where $n$ stands for the number of nodes in the graph, and $p$ stands for the independent edge connection probability in the graph. As shown in Figure 1, these three representations can have overlapping to mutually contribute to each other Fu and He (2022).
@@ 1.3 Organization of this Paper
Graph law is the study of investigating the statistical properties of graphs. In this survey, we introduce the graph laws studies in the macroscopic view and microscopic view, plus multiple angles like low-order and high-order connections, static and dynamic graphs, as shown in Table 1.
  • Macroscopic and Microscopic Views. The macroscopic graph laws describe the graph properties in a global view, like how the total degree (or eigenvalues) distribution of the entire graph looks like Leskovec et al. (2005); while the microscopic laws try to focus on the individual behavior and investigate their behaviors as part in the entire graph Leskovec et al. (2008).
  • Low-Order and High-Order Connections. Most graph laws are based on the node-level connections (i.e., low-order connections), while some graph law investigations are based on the group activities (high-order connections), i.e., motif in Paranjape et al. (2017); Zeno et al. (2020), hyperedge in Do et al. (2020); Kook et al. (2020), and simplex in Benson et al. (2018); Comrie and Kleinberg (2021).
  • Static and Dynamic Graphs. Compared with static graphs, dynamic graphs allow the graph components like topology structure and node attributes to evolve over time. Correspondingly, some graph laws study how the graph parameters change over time and their temporal relations. Note that, in some graph research, the dynamics are created by the algorithms, like adding virtual nodes to preserve graph representations Abraham et al. (2007); Liu et al. (2022); this kind of research is beyond the discussion of this paper, and we focus on the graph law with natural time.
Furthermore, in Section 4, we survey different real-world applications that would benefit from the guidance of graph laws. In Section 5, we introduce different observation spaces and newly proposed parameters, whose corresponding laws are not fully discovered yet. Also, in Section 6, we review the related work of graph law surveys and compare the differences. Finally, in Section 7, we conclude the survey with current challenges and future directions.
@@
Graphs Densification Law Leskovec et al. (2005) Density degree $\alpha$ Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\in[1,2]$, $e(t)$ is # edges at $t$
@@
Shrinking Law Leskovec et al. (2005) Effective diameter $d$ Macro Low Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
@@
Motif Differing Law(1) Paranjape et al. (2017) Numbers of similar motifs $n$ Macro High Dynamic $n_{1}\neq n_{2}$ for different domains
@@
Motif Differing Law(2) Paranjape et al. (2017) Motif occurring timestamp $t$ Macro High Dynamic $t_{1}\neq t_{2}$ for different motifs
@@
Egonet Differing Law Benson et al. (2018) Features of Egonets $X$ Macro High Static $X_{1}\neq X_{2}$ for different domains
@@
Simplicial Closure Law Benson et al. (2018) Simplicial closure probability $p$ Macro High Static $p$ increases with additional edges or tie strength
@@
Spectral Power Law(1) Eikmeier and Gleich (2017) Degree, SVD, eigen distributions Macro High Static These distributions usually follow power-law
@@
Spectral Power Law(2) Eikmeier and Gleich (2017) Degree, SVD, eigen distributions Macro High Static If one follow power-law, usually others follow
@@
Edge Attachment Law(1) Leskovec et al. (2008) Node degree $d$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)\propto d$ for node with degree $d$
@@
Edge Attachment Law(2) Leskovec et al. (2008) Node age $a(u)$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)$ seems to be non-decreasing with $a(u)$
@@
Triangle Closure Law(1) Huang et al. (2018) Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{3}\Rightarrow$ unlikely $e_{1}/e_{2}$ will be weakened
@@
Triangle Closure Law(2) Huang et al. (2018) Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{1}/e_{2}\Rightarrow$ unlikely they will be weakened
@@
Local Closure Law Yin et al. (2019) Local closure coefficient $H(u)$ Micro Low Static Please refer to Section 5 for details
@@
Spectral Density Law Dong et al. (2019) Density of states $\mu(\lambda)$ Macro High Static Please refer to Section 5 for details
@@
Motif Activity Law(1) Zeno et al. (2020) Motif type Micro High Dynamic Motifs do not transit from one type to another
@@
Motif Activity Law(2) Zeno et al. (2020) Motif re-appear rate Micro High Dynamic Motifs re-appear with configured rates
@@
Hypergraphs Degree Distribution Law Do et al. (2020) Node degree, edge link probability Macro High Dynamic High-degree nodes are likely to form new links
@@
SVD Distribution Law Do et al. (2020) Singular value distribution Macro High Static Singular value distribution usually heavy-tailed
@@
Diminishing Overlaps Kook et al. (2020) density of interactions $DoI(\mathcal{H}(t))$ Macro High Dynamic Overall hyperedge overlaps decrease over time
@@
Densification Law Kook et al. (2020) Density degree $\alpha$ Macro High Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$, $e(t)$ is # hyperedges at $t$
@@
Shrinking Law Kook et al. (2020) Hypergraph effective diameter $d$ Macro High Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
@@
Edge Interacting Law Comrie and Kleinberg (2021) Edge interacting rate Micro High Dynamic Temporally adjacent interactions highly similar
@@
Heterographs Densification Law Wang et al. (2019) Density degree $\alpha$, # meta-path Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$ for some meta-path
@@
Non-densification Law Wang et al. (2019) Density degree $\alpha$, # meta-path Macro Low Dynamic Maybe, for some meta-path, $e(t)\not\propto n(t)^{\alpha}$
@@ 2 Macroscopic Graph Laws
Several classical theories model the growth of graphs, for example, Barabasi-Albert model Barabási and Albert (1999); Barabâsi et al. (2002) assumes that the graphs follow the uniform growth pattern in terms of the number of nodes, and the Bass model Mahajan et al. (1990) and the Susceptible-Infected model Anderson and May (1992) follow the Sigmoid growth (more random graph models can be founded in Drobyshevskiy and Turdakov (2020)). However, these pre-defined graph growths have been tested that they could not handle the complex real-world network growth patterns very well Kwak et al. (2010); Zang et al. (2018). To this end, researchers begin to fit the graph growth on real-world networks directly, to discover graph laws.
@@ 2.1 Low-Order Macroscopic Laws
Based on fitting nine real-world temporal graphs from four different domains, the authors in Leskovec et al. (2005) found two temporal graph laws, called (1) Densification Laws and (2) Shrinking Diameters, respectively. First, the densification law states as follows.
@@
where $e(t)$ denotes the number of edges at time $t$, $n(t)$ denotes the number of nodes at time $t$, $\alpha\in[1,2]$ is an exponent representing the density degree. The second law, shrinking diameters, states that the effective diameter is decreasing as the network grows, in most cases. Here, the diameter means the node-pair shortest distance, and the effective diameter of the graph means the minimum distance $d$ such that approximately 90% of all connected pairs are reachable by a path of length at most $d$. Later, in Zang et al. (2018), the densification law gets in-depth confirmed on four different real social networks, the research shows that the number of nodes and number of edges both grown exponentially with time, i.e., following the power-law distribution.